# Blind Inverse Problem Solving Made Easy by Text-to-Image Latent Diffusion

### v1

@@ Appendix D Limitations and Future Works
Similar to many training-free diffusion posterior sampling algorithms [8, 27, 9, 14], our method is also sensitive to hyper-parameter tuning. We have provided details about our hyperparameter choices in the previous sections, and we will release our code in a public repository upon publication of this paper.
Another notable drawback of our method is that LADiBI requires significantly longer inference time in comparison to the best performing baseline (i.e. GibbsDDRM [27]) in order to obtain high quality restorations. With the general operator initialization, our algorithm can take around 5 minutes to complete on a single NVIDIA A6000 GPU while GibbsDDRM only takes around 30 seconds. Although this is justifiable by the larger optimization space that we operate on, investigating on how to reduce the inference time requirement is an interesting and critical next step for our work.
@@
Appendix E Broader Impact Statement

### v2

@@ Appendix E Limitations and Future Works
Similar to many training-free diffusion posterior sampling algorithms [Chung et al., 2023a, Murata et al., 2023, Chung et al., 2023b, He et al., 2024], our method is also sensitive to hyper-parameter tuning. We have provided details about our hyperparameter choices in the previous sections, and we will release our code in a public repository upon publication of this paper.
In addition, while generally simple, we do require the users to infer appropriate prompts from the measurement. An interesting future work direction can include automated prompt tuning similar to the method proposed in Chung et al. [2023c].
Another notable drawback of our method is that LADiBI requires significantly longer inference time in comparison to the best performing baseline (i.e. GibbsDDRM [Murata et al., 2023]) in order to obtain high quality restorations. With the general operator initialization, our algorithm can take around 5 minutes to complete on a single NVIDIA A6000 GPU while GibbsDDRM only takes around 30 seconds. Although this is justifiable by the larger optimization space that we operate on, investigating on how to reduce the inference time requirement is an interesting and critical next step for our work.
@@
Appendix F Impact Statement